OpenStudy (anonymous):

Use basic identities to simplify the expression. cosecant of theta times cotangent of theta divided by secant of theta.

OpenStudy (anonymous):

\[ \frac{ \csc \theta \cot \theta }{ \sec \theta }\]

OpenStudy (anonymous):

I already have all the identities written, I just don't know how to apply them to the equation.

OpenStudy (anonymous):

I'm aware that csc is 1/sin and cot is sin/cos while sec is 1/cos but there should be something that goes between all of that to give me the answer.

OpenStudy (anonymous):

Sorry, cot is cos/sin.